# Msmsfnet: a multi-stream and multi-scale fusion net for edge detection

### v1

@@ IV Conclusions
In this paper we focus on the task of edge detection in natural images. Unlike previous works that use pretrained ImageNet models, in this work all models are trained from scratch and evaluated under the same experimental settings. By doing so, we ensure the fairness to compare the detection capability of different models. In the meantime, the experimental settings in this work facilitate the design of new network architectures for edge detection, as pretraining a model on the ImageNet dataset is not pratical for many researchers due to the limited computation resources. Besides, we propose a new model for edge detection, which adopts a multi-branch architecture to enhance its multi-scale representation capability and utilizes spatial asymmetric convolutions so as to increase its accuracy by developing a very deep model. The proposed msmsfnet achieves superior performance than state-of-the-art methods in three publicly available datasets under the same experimental settings, which shows the efficiency of the proposed method.
Source code
The source code of the method will be released upon the acceptance of the paper .
@@
This work is supported in part by the National Natural Science Foundation of China (Grant No. 42374018), in part by the Shenzhen Scientific Research and the Development Funding Program (Grant No. RCYX20210706092140076), in part by the Swift Fund Fintech Funding and the Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ, Grant No. GML-KF-22-15), in part by the Shenzhen Science and Technology Program (Grant No. JCYJ20220531101409021), and in part by the National Natural Science Foundation of China (Grant No. 62371348).

### v3

@@ 5 Conclusions
In this paper we focus on the task of edge detection in both natural images and SAR images. Unlike previous works that use pre-trained ImageNet models, we study the performance achievable by existing methods when the pre-trained weights are not used. The motivation behind this work is two-fold: first, the use of pre-trained weights in existing methods limits the design space of network architecture for edge detection, as the proposed methods have to be built upon existing well-trained ImageNet models; second, the strategy to pre-train a model using a large dataset and then fine-tune the model using a small dataset, which is commonly used in the field of computer vision, is not always possible, as explained and demonstrated in the experiments for edge detection in SAR images. The experimental settings in this work facilitate the design of new network architectures for edge detection, as pre-training a model on the ImageNet dataset is not practical for many researchers due to limited computation resources. Furthermore, we propose a new model for edge detection, which adopts a multi-branch architecture to enhance its multi-scale representation capability and utilizes spatial asymmetric convolutions so as to increase its accuracy by developing a very deep model. The proposed msmsfnet achieves superior performance than state-of-the-art methods when all models are trained from scratch. When the pre-trained weights are used for parameter initialization, the performance of the proposed msmsfnet improves further and is also competitive on the BSDS500 dataset, which shows the efficiency of the proposed method.
@@
This work is supported in part by the National Natural Science Foundation of China (Grant No. 42374018 and Grant No. 42304012), in part by the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2025B1515020092 and Grant No. 2022A1515110730), in part by the Shenzhen Science and Technology Program (Grant No. KCXFZ20240903093000002 and Grant No. JCYJ20220531101409021). and in part by the National Natural Science Foundation of China (Grant No. 62371348).